OpenStudy (anonymous):

l (7/9x) +4 l +6=6 solve the absolute value

OpenStudy (dumbcow):

well the 6's cancel |(7/9x) +4| = 0 since -0 is same as +0, absolute value is irrelevant in this case (7/9x) +4 = 0 7/9x = -4 7 = -36x -7/36 = x